OpenStudy (anonymous):

@phi HOW DIFFERENTIATE THIS ONE: y=9^x-5^x

OpenStudy (phi):

the "easy" way to remember is rewrite 9 as e^ln(9) and raise to the x power: e^(x*ln9) the derivative is ln9* e^(x*ln9)= ln9*9^x

OpenStudy (phi):

but I think people just memorize d/dx n^x is ln(n)* n^x

OpenStudy (anonymous):

(a^x) ' = a^x * ln a

OpenStudy (anonymous):

oh. I got it

OpenStudy (anonymous):

THANX GUYS

OpenStudy (anonymous):

:)